# The Labyrinth of Links: Navigating the Associative Maze of Multi-modal LLMs

### v1

@@ Abstract
Multi-modal Large Language Models (MLLMs) have exhibited impressive capability. However, recently many deficiencies of MLLMs have been found compared to human intelligence, e.g., hallucination. To drive the MLLMs study, the community dedicated efforts to building larger benchmarks with complex tasks. In this paper, we propose benchmarking an essential but usually overlooked intelligence: association, a human’s basic capability to link observation and prior practice memory. To comprehensively investigate MLLM’s performance on the association, we formulate the association task and devise a standard benchmark based on adjective and verb semantic concepts. Instead of costly data annotation and curation, we propose a convenient annotation-free construction method transforming the general dataset for our association tasks. Simultaneously, we devise a rigorous data refinement process to eliminate confusion in the raw dataset. Building on this database, we establish three levels of association tasks: single-step, synchronous, and asynchronous associations. Moreover, we conduct a comprehensive investigation into the MLLMs’ zero-shot association capabilities, addressing multiple dimensions, including three distinct memory strategies, both open-source and closed-source MLLMs, cutting-edge Mixture-of-Experts (MoE) models, and the involvement of human experts. Our systematic investigation shows that current open-source MLLMs consistently exhibit poor capability in our association tasks, even the currently state-of-the-art GPT-4V(vision) also has a significant gap compared to humans. We believe our benchmark would pave the way for future MLLM studies. Our data and code are available at: https://mvig-rhos.com/llm_inception.
@@ 1 Introduction
In this paper, we aim to devise a standard benchmark to evaluate MLLM’s capability in association tasks. As shown in Figure 1a, we always connect observations with previous practice memory. This association is usually the semantic concept shared between different objects, representing common features in general dataset samples. For instance, the “painted airplane” and “painted bus” share the same attribute of “painted”, which can be seen as an association chain in the association task. Next, it is natural to ask can we design a general method to build an association benchmark based on existing datasets without too much extra effort? To this end, we propose an annotation-free association construction method, which easily transfers the general dataset for the association task. Thus, as in Figure 1b, we devise a standard association benchmark using the object’s concepts as an association chain. Given the benchmark, we comprehensively investigate MLLMs’ capabilities in association tasks across multiple memory strategies, multiple open-source and closed-source MLLMs, as well as cutting-edge Mixture-of-Experts (MoE) and human expert testers.
@@ 2 Related works
Multi-Modal Large Language Model. There has been a surge of MLLMs in the deep learning community, which effectively use off-the-shell LLMs to support multi-modal inputs and demonstrate promise for zero-shot generalization. For instance, LLaVA Liu et al. (2023a) InstructBLIP Dai et al. (2023), InternLM-XComposer Dong et al. (2024), Qwen-VL Bai et al. (2023a), MiniGPT-4 Zhu et al. (2023), and mPLUG-Owl2 Ye et al. (2023) make leading in last year according to their powerful visual perception capabilities. More recently, these models Wang et al. (2024b); Li et al. (2024a); Ye et al. (2024) are updated to improve their ability in multi-images or video input. In this work, we mainly evaluated the leading open-source models Qwen2-VL Wang et al. (2024b), mPLUG-Owl3 Ye et al. (2024), and LLaVA-OneVision Li et al. (2024a).
Tuning-Free Engineering. The development of MLLMs has led to a significant improvement in in-context learning Doveh et al. (2024); Wies et al. (2024). Additionally, the use of visual and language prompts has been shown to enhance the performance of MLLMs. Research has indicated that visual cues such as listed numbers Yan et al. (2024) and shapes Shtedritski et al. (2023); Mani et al. (2020) can be effectively understood by MLLMs. Similarly, carefully designed language prompts Liu et al. (2023b), Chain-of-Thought (CoT) Wei et al. (2024), tree-of-mixed-thought (ToMT) Hu et al. (2023), have shown promising results. Common knowledge Shao et al. (2023); Liu et al. (2024) can be seen as an instruction Prior knowledge about how to decisions on a task. As a result, we chose to employ the general prompt-engineer one-shot, CoT, and Common knowledge.
@@ 3 Annotation-free Association Construction
For a general dataset with $N$ samples, it can be formed as $\{(x_{i},y_{i})\mid x_{i}\in X,y_{i}\in\{c_{1},\dots,c_{n}\},i=1,2,\dots,N\}$. Let $x_{i}=(x_{i}^{m_{1}},\ldots,x_{i}^{m_{k}})$ be a sample with $k$ modalities, more specifically, for a task with individual modality, the sample is described as $x_{i}=(x_{i}^{m_{1}},)$. $y_{i}$ is the annotation for sample $x_{i}$, which may existed in various granularities. These annotations typically indicate a semantic concept, with object categories as nouns, actions as verbs, as well as attributes, and affordance as adjectives.
Given the definition, labels represent a subset of concepts present in the given sample. The core of human association, on the other hand, involves identifying the overlapping concepts between newly acquired observations and prior practice memory. Hence, it is intuitive to use annotations from the raw dataset to construct the association chain, which reflects the common concepts shared by two samples, such as the presence of the “shoot ball” action in both samples. In the following, we first introduce the method for generating association chains. We then introduce a method that deduces the evidence of the association to simulate the thought behind the human, which involves predicting the common concepts among the samples that constitute the association.
@@
For a general dataset $\{(x_{i},y_{i})\mid x_{i}\in X,y_{i}\in\{c_{1},\dots,c_{n}\},i=1,2,\dots,N\}$, we randomly select two samples to form a new sample pair for association tasks. If the selected samples have identical labels or share at least one common label, we assign a corresponding label of $1$; otherwise, we assign a label of 00. Hence, we get the new association dataset as:
@@
In this way, for each sample $x_{i}$ in the original dataset, we construct a positive association set with $K$ positive samples, $\bm{x}_{i}^{+}=[p^{1}_{i},\ldots,p^{K}_{i}]$, where at least one potential association concept exists. Additionally, we devise another negative association set with $L$ samples, $\bm{x}_{i}^{-}=[q_{i}^{1},\ldots,q_{i}^{L}]$, in which no association chain is present.
@@
It is crucial to deduce the evidence in the association task, which simulates the human reflection process. This evidence then guides and directs the decision-making in the subsequent steps. To realize this, we collected the full set of shared concepts within all possible positive set $\bm{x}_{i}^{+}$ as $\hat{C}=\bigcup_{i=1}^{N}\left\{z_{ij}\mid z_{ij}=y_{i}\cap y_{j},1\leq i<j\leq N\right\}.$ In this setting, the dataset deducting the evidence of the association chain is depicted as:
@@
Based on the above definition, we can easily construct the association task. First, we randomly take one sample, $x_{\text{query}}$, as the initial query image. For $t$-th step, We then randomly choose candidate samples $p_{i}^{t}$ and $q_{i}^{t}$ from the positive $\bm{x}^{+}_{\text{query}}$ and negative $\bm{x}^{-}_{\text{query}}$ association sets, respectively. Next, we use the MLLMs to determine which candidate samples share common concepts with the query image, the output of $t$-th step described as $o_{t}$. If the model makes the correct decision, $o_{t}=p_{i}^{t}$, we set the correctly predicted sample as the new query image and repeatedly select candidate samples and make decisions. Otherwise, we calculate the maximum association steps and exit. The above description can be depicted as:
@@
where $x_{\text{query}}\in\{x_{1},\ldots,x_{N}\}$ and $\mathcal{F}$ represents the forward computation process of MLLMs. step is the computation of the individual association step. Furthermore, after each association step, we will deduce the evidence of the association chain, which will integrated into subsequent associations.
@@ 4 Association Benchmark
In this section, we introduce a benchmark based on the semantic concepts of adjectives and verbs, i.e., object attributes and affordances, and human actions. Specifically, we utilize the annotation-free construction method proposed in Section 3 on the Object Concept Learning (OCL) Li et al. (2023) to generate an attribute and affordance association datasets, and on the Pangea Li et al. (2024b) to generate action association dataset  [^1] To further demonstrate the capability on verb concept, we further implemented on action HMDB Kuehne et al. (2011) datasets. The results are included in the section D of the supplementary material.. In the following, we comprehensively investigate the association ability from the single-step association, synchronous association, and asynchronous association settings as shown in Figure 2.
@@ 4.2 Data Refinement for Association Task
We implemented a three-step strategy to ensure data quality, including an Image resolution filter, MLLM verification, and Human expert evaluation. Specifically, the Image resolution filter screened out all images with less than $50,000$ pixels to ensure superior visual quality. The MLLM verification  [^2] We compare the performance with and without the MLLM verification in the section E of supplementary. takes a question-answer strategy with OpenAI’s GPT4-V Achiam et al. (2023) and Google’s Gemini-1.5-Flash Reid et al. (2024) to ensure each annotation of raw data exists in the image. Then, the human expert evaluation is conducted through our custom-designed interface, enabling testers to complete the association task and eliminate low-quality samples or those with ethical concerns  [^3] The human expert test interface refers to Figure 7 in the supplementary..
Our benchmarks are implemented in adjective concepts and verb concepts, which include attribute and affordance in OCL Li et al. (2023) and action in Pangea Li et al. (2024b). In the OCL dataset, we selected eight attributes with good perception performance, such as “metal, ripe, fresh, natural, cooked, painted, rusty, furry”, and eight affordances “sit, imprint, push, carry, cut, clean, open, break”. In addition, we selected eight actions of “run, hit, drive, dress, cooking, build, shake, cut”.
@@ 5.1.1 Implementation Details
We systematically conduct experiments that include concept perception, single-step, synchronous, and asynchronous association. Specifically, we implement the task as a multi-choice setting in all experiments, taking one query image and two candidate images with one correct as input and output correct image index. Based on this, we first devise preliminary concept perceptions that involve popular prompt engineering skills on open-source MLLMs  [^4] For the detailed description of the setting of concept perception, please refer to section B. to investigate perception capabilities in attribute concept. Then, we convert to evaluate MLLM’s association capabilities that make decisions based on current observation and prior practice memory, i.e., input with additional content of previous practice. We develop single, synchronous, and asynchronous associations according to fixed or dynamic memory. The single-step association means the model decides with observation and correct prior practice. Meanwhile, the synchronous and asynchronous association set model at a dynamic task that iteratively arrives at a decision and then deducts the underlying evidence, which means the memory may have wrong information for the next judge.
@@
We utilize three open-source MLLMs in preliminary concept perception: QWen-VL Bai et al. (2023b), LLaVA-NeXT-7B Liu et al. (2024), and LLaVA-NeXT-13B. For formal association, we utilize three new-versions MLLMs that break through the MLLM’s capabilities in multi-images: LLaVA-OneVision Li et al. (2024a), QWen2-VL Wang et al. (2024b), and mPLUG-Owl3 Ye et al. (2024). Besides, we evaluate the performance of Mixture-of-Experts (MoE) that combined three open-source MLLMs. In experiments, open-source MLLM is run on a single NVIDIA A100 80G GPU. Apart from open-source MLLMs, we include the evaluation of the closed-source MLLMs of GPT4-V Achiam et al. (2023) and Gemini-1.5-Flash Reid et al. (2024). Simultaneously, we involve the results of three human experts to demonstrate the gap between MLLM with human intelligence.
@@ 5.2 Relationship between Perception and Association
In this subsection, we investigate the MLLM’s ability in attribute concept perception that is implemented on the OCL Li et al. (2023). We devise multiple task complexity, including individual perception and combination perception. The detailed description and results are shown in section B in the supplementary. These results indicate that while popular prompt engineering skills improve performance in specific cases, i.e., prompt engineering skills compared to Normal, they still exhibit limited capabilities in the widely-used object understanding benchmark. For instance, in the individual perception, the highest perception success ratio of $0.690$ reflects an improvement of $0.190$ over the random baseline of $0.5$. Furthermore, the combination perception attains the success ratio of $0.542$, improving $0.292$ over the random baseline of $0.25$.
@@
Type Concept Memory Models (Success Ratio)
LLaVa-OneVision QWen2-VL mPLUG-Owl3 Gemini-1.5-Flash Avg.
Association Success Ratio Attribute (Adjective) NoM $75.52$ $76.36$ $54.05$ $75.40$ $70.33$
StructM $78.41$ $77.74$ $72.67$ $87.30$ $79.03$
NLM $80.40$ $86.63$ $73.33$ 88.39 $82.19$
ChainM $77.73$ $73.55$ $70.70$ $83.78$ $76.44$
Affordance (Adjective) NoM $70.61$ $75.10$ $53.80$ $77.78$ $69.32$
StructM $73.39$ $75.39$ $73.02$ $\underline{84.45}$ $76.56$
NLM $76.13$ $79.37$ $67.79$ 85.40 $77.17$
ChainM $74.38$ $82.67$ $68.46$ $80.02$ $76.38$
Action (Verb) NoM $75.74$ $78.43$ $57.94$ $84.21$ $74.08$
StructM $75.44$ $82.10$ $73.92$ $\underline{88.10}$ $79.89$
NLM $78.66$ $88.01$ $70.04$ 89.58 $81.57$
ChainM $76.92$ $85.59$ $69.50$ $87.58$ $79.90$
Deduction Success Ratio Attribute (Adjective) $49.38$ $58.11$ $61.21$ $65.82$ $58.63$
Affordance (Adjective) $21.07$ $15.98$ $33.30$ $33.08$ $25.86$
Action (Verb) $46.61$ $49.99$ $42.89$ $55.69$ $48.80$

### v2

@@ Abstract
Multi-modal Large Language Models (MLLMs) have exhibited impressive capability. However, recently many deficiencies of MLLMs have been found compared to human intelligence, e.g., hallucination. To drive the MLLMs study, the community dedicated efforts to building larger benchmarks with complex tasks. In this paper, we propose benchmarking an essential but usually overlooked intelligence: association, a human’s basic capability to link observation and prior practice memory. To comprehensively investigate MLLM’s association performance, we formulate the association task and devise a standard benchmark based on adjective and verb semantic concepts. Instead of costly data annotation and curation, we propose a convenient annotation-free construction method transforming the general dataset for our association tasks. Simultaneously, we devise a rigorous data refinement process to eliminate confusion in the raw dataset. Building on this database, we establish three levels of association tasks: single-step, synchronous, and asynchronous associations. Moreover, we conduct a comprehensive investigation into the MLLMs’ zero-shot association capabilities, addressing multiple dimensions, including three distinct memory strategies, both open-source and closed-source MLLMs, cutting-edge Mixture-of-Experts (MoE) models, and the involvement of human experts. Our systematic investigation shows that current open-source MLLMs consistently exhibit poor capability in our association tasks, even the currently state-of-the-art GPT-4V(vision) also has a significant gap compared to humans. We believe our benchmark would pave the way for future MLLM studies. Our data and code are available at: https://mvig-rhos.com/llm_inception.
@@ 1 Introduction
In this paper, we aim to devise a standard benchmark to evaluate MLLM’s capability in association tasks. As shown in Figure 1a, we always connect observations with previous practice memory. This association is usually the semantic concept shared between different objects, representing common features in general dataset samples. For instance, the “painted airplane” and “painted bus” share the same attribute of “painted”, which can be seen as an association chain in the association task. Next, it is natural to ask can we design a general method to build an association benchmark based on existing datasets without too much extra effort? To this end, we propose an annotation-free association construction method, which easily transfers the general dataset for the association task. As in Figure 1b, we devise a standard association benchmark using the object’s concepts as an association chain. Given the benchmark, we comprehensively investigate MLLMs’ capabilities in association tasks across multiple memory strategies, multiple open-source and closed-source MLLMs, as well as cutting-edge Mixture-of-Experts (MoE) and human expert testers.
@@ 2 Related works
Multi-Modal Large Language Model. There has been a surge of MLLMs in the deep learning community, which use off-the-shell LLMs to support multi-modal inputs and demonstrate promise for zero-shot generalization. For instance, LLaVA Liu et al. (2023a) InstructBLIP Dai et al. (2023), InternLM-XComposer Dong et al. (2024), Qwen-VL Bai et al. (2023a), MiniGPT-4 Zhu et al. (2023), and mPLUG-Owl2 Ye et al. (2023) make leading in last year according to their powerful visual perception capabilities. More recently, these models Wang et al. (2024b); Li et al. (2024a); Ye et al. (2024) are updated to improve their ability in multi-images or video input. In this work, we mainly evaluated the leading open-source models Qwen2-VL Wang et al. (2024b), mPLUG-Owl3 Ye et al. (2024), and LLaVA-OneVision Li et al. (2024a).
Tuning-Free Engineering. The development of MLLMs has led to a significant improvement in in-context learning Doveh et al. (2024); Wies et al. (2024). Additionally, the use of visual and language prompts has been shown to enhance the performance of MLLMs. Research has indicated that visual cues such as listed numbers Yan et al. (2024) and shapes Shtedritski et al. (2023); Mani et al. (2020) can be effectively understood by MLLMs. Similarly, carefully designed language prompts Liu et al. (2023b), Chain-of-Thought (CoT) Wei et al. (2024), tree-of-mixed-thought (ToMT) Hu et al. (2023), have shown promising results. Common knowledge Shao et al. (2023); Liu et al. (2024a) can be seen as an instruction of Prior knowledge about how to make decisions on a task. As a result, we chose to employ the general prompt-engineer one-shot, CoT, and Common knowledge.
@@ 3 Annotation-free Association Construction
For a general dataset with $N$ samples, it can be formed as $\{(x_{i},y_{i})\mid x_{i}\in X,y_{i}\in\{c_{1},\dots,c_{n}\},i=1,2,\dots,N\}$. Let $x_{i}=(x_{i}^{m_{1}},\ldots,x_{i}^{m_{k}})$ be a sample with $k$ modalities, more specifically, for a task with individual modality, the sample is described as $x_{i}=(x_{i}^{m_{1}},)$. $y_{i}$ is the annotation for sample $x_{i}$, which may existed in various granularities. These annotations indicate a semantic concept, with object categories as nouns, actions as verbs, as well as attributes, and affordance as adjectives.
Given this definition, labels represent a subset of link concepts present in the given sample. The core of human association, on the other hand, involves identifying the overlapping concepts between newly acquired observations and prior practice memory. Hence, it is intuitive to use annotations from the raw dataset to construct the association chain, which reflects the common concepts shared by two samples, such as the presence of the “shoot ball” action in both samples. In the following, we first generate possible chains for the association task, then create the ground-truth labyrinth to reason the chain behind the association.
@@
Association connects objects by any potential links. We defined associations existing when any pre-defined semantic concept appears between objects. To this end, for a general dataset $\{(x_{i},y_{i})\mid x_{i}\in X,y_{i}\in\{c_{1},\dots,c_{n}\},i=1,2,\dots,N\}$, we randomly select two samples to form a new sample pair for association. If the selected samples have identical labels or share at least one common label, we assign a corresponding label of $1$; otherwise, we assign a label of 00. Hence, we get the new association dataset as:
@@
In this way, for each sample $x_{i}$ in the original dataset, we construct a positive association set with $K$ positive samples, $\bm{x}_{i}^{+}=[p^{1}_{i},\ldots,p^{K}_{i}]$, where at least one potential association concept exists. Additionally, we devise another negative association set with $L$ samples, $\bm{x}_{i}^{-}=[q_{i}^{1},\ldots,q_{i}^{L}]$, in which no predefined association chain present.
@@
Associations use implicit links to connect objects, which make correct decisions with unspecified links. To uncover the labyrinth, we devise deduction steps that reason the association links after association. This evidence will fused into memory and influence subsequent steps. To realize this, we collected the full set of shared concepts within all possible positive set $\bm{x}_{i}^{+}$ as $\hat{C}=\bigcup_{i=1}^{N}\left\{z_{ij}\mid z_{ij}=y_{i}\cap y_{j},1\leq i<j\leq N\right\}.$ In this setting, the dataset deducting the evidence of the association chain is depicted as:
@@
Based on the above definition, we can easily construct the association task. First, we randomly select one sample, $x_{\text{query}}$, as the initial query image. Then, for each step ${t}$, We randomly choose candidate samples $p_{i}^{t}$ and $q_{i}^{t}$ from the positive $\bm{x}^{+}_{\text{query}}$ and negative $\bm{x}^{-}_{\text{query}}$ sets, respectively. Next, we deliver new observations and previous practice memory into MLLMs determining which candidate shares common concepts with the query, the output of $t$-th step described as $o_{t}$. If the model makes the correct decision, $o_{t}=p_{i}^{t}$, the query and correctly predicted images will be used to deduct the chain of association that updates memory. Then convert to the next association step, the correctly predicted samples taken as the query image and repeatedly select candidate samples to conduct association and deduction progress. If make an error, we calculate the maximum association steps and exit. The above description can be depicted as:
@@
where $x_{\text{query}}\in\{x_{1},\ldots,x_{N}\}$ and $\mathcal{F}$ represents the forward computation of MLLMs. step is the computation of combined steps of association and deduction.
@@ 4 Association Benchmark
In this section, we introduce a benchmark based on the semantic concepts of adjectives and verbs, i.e., object attributes and affordances, and human actions. Specifically, we utilize the annotation-free construction method proposed in Section 3 on the Object Concept Learning (OCL) Li et al. (2023) to generate an attribute and affordance association datasets, and on the Pangea Li et al. (2024c) to generate action association dataset  [^1] To further demonstrate the capability on verb concept, we further implemented on action HMDB Kuehne et al. (2011) datasets. The results are included in the section F of the supplementary material.. In the following, we comprehensively investigate the association ability from the single-step association, synchronous association, and asynchronous association settings as shown in Figure 2.
@@ 4.2 Data Refinement for Association Task
We implemented a three-step strategy to ensure data quality, including an Image resolution filter, MLLM verification, and Human expert evaluation. Specifically, the Image resolution filter screened out all images with less than $50,000$ pixels to ensure superior visual quality. The MLLM verification  [^2] We compare the performance with and without the MLLM verification in the section G of supplementary. takes a question-answer strategy with OpenAI’s GPT4-V Achiam et al. (2023) and Google’s Gemini-1.5-Flash Reid et al. (2024) to ensure each annotation of raw data exists in the image. Then, the human expert evaluation is conducted through our custom-designed interface, enabling testers to complete the association task and eliminate low-quality samples or those with ethical concerns  [^3] The detailed visualization and description refer to Subsection A.1 in the supplementary..
Our benchmarks are implemented in adjective concepts and verb concepts, which include attribute and affordance in OCL Li et al. (2023) and action in Pangea Li et al. (2024c). In the OCL dataset, we selected eight attributes with good perception performance, such as “metal, ripe, fresh, natural, cooked, painted, rusty, furry”, and eight affordances “sit, imprint, push, carry, cut, clean, open, break”. In addition, we selected eight actions of “run, hit, drive, dress, cooking, build, shake, cut”.
@@ 5.1.1 Implementation Details
We systematically conduct experiments that include concept perception, single-step, synchronous, and asynchronous association. Specifically, we implement the task as a multi-choice setting in all experiments, taking one query image and two candidate images with one correct as input and output correct image index. Based on this, we first devise preliminary concept perceptions that involve popular prompt engineering skills on open-source MLLMs  [^4] For the detailed description of the setting of concept perception, please refer to section D. to investigate perception capabilities in attribute concept. Then, we convert to evaluate MLLM’s association capabilities that make decisions based on current observation and prior practice memory, i.e., input with additional content of previous practice. We develop single, synchronous, and asynchronous associations according to fixed or dynamic memory. The single-step association means the model decides with observation and correct prior practice. Meanwhile, the synchronous and asynchronous association set model at a dynamic task that iteratively arrives at a decision and then deducts the underlying evidence, which means the memory may have wrong information for the next judge.
@@
We utilize three open-source MLLMs in preliminary concept perception: QWen-VL Bai et al. (2023b), LLaVA-NeXT-7B Liu et al. (2024a), and LLaVA-NeXT-13B. For formal association, we utilize three new-versions MLLMs that break through the MLLM’s capabilities in multi-images: LLaVA-OneVision Li et al. (2024a), QWen2-VL Wang et al. (2024b), and mPLUG-Owl3 Ye et al. (2024). Besides, we evaluate the performance of Mixture-of-Experts (MoE) that combined three open-source MLLMs. In experiments, open-source MLLM is run on a single NVIDIA A100 80G GPU. Apart from open-source MLLMs, we include the evaluation of the closed-source MLLMs of GPT4-V Achiam et al. (2023) and Gemini-1.5-Flash Reid et al. (2024). Simultaneously, we involve the results of three human experts to demonstrate the gap between MLLM with human intelligence.
@@ 5.2 Relationship between Perception and Association
In this subsection, we investigate the MLLM’s ability in attribute concept perception that is implemented on the OCL Li et al. (2023). We devise multiple task complexity, including individual perception and combination perception. The detailed description and results are shown in section D in the supplementary. These results indicate that while popular prompt engineering skills improve performance in specific cases, i.e., prompt engineering skills compared to Normal, they still exhibit limited capabilities in the widely-used object understanding benchmark. For instance, in the individual perception, the highest perception success ratio of $0.690$ reflects an improvement of $0.190$ over the random baseline of $0.5$. Furthermore, the combination perception attains the success ratio of $0.542$, improving $0.292$ over the random baseline of $0.25$.